### v1

@@ Abstract
Current studies have exposed the risk of Large Language Models (LLMs) generating harmful content by jailbreak attacks. However, they overlook that the direct generation of harmful content from scratch is more difficult than inducing LLM to calibrate benign content into harmful forms. In our study, we introduce a novel attack framework that exploits AdVersArial meTAphoR (AVATAR) to induce the LLM to calibrate malicious metaphors for jailbreaking. Specifically, to answer harmful queries, AVATAR adaptively identifies a set of benign but logically related metaphors as the initial seed. Then, driven by these metaphors, the target LLM is induced to reason and calibrate about the metaphorical content, thus jailbroken by either directly outputting harmful responses or calibrating residuals between metaphorical and professional harmful content. Experimental results demonstrate that AVATAR can effectively and transferable jailbreak LLMs and achieve a state-of-the-art attack success rate across multiple advanced LLMs.
@@
Yu Yan1,2, Sheng Sun1, Zenghao Duan1,2, Teli Liu3, Min Liu5, Zhiyi Yin1, Qi Li4, Jingyu Lei4, 1Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China 2University of Chinese Academy of Sciences, Beijing, China 3People’s Public Security University of China, Beijing, China 4Tsinghua University, Beijing, China 5Zhongguancun Laboratory, Beijing, China {yanyu24z, sunsheng, duanzenghao24s, liumin}@ict.ac.cn, {qli01, leijy}@tsinghua.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) Yi et al. (2024); Achiam et al. (2023) have become increasingly prevalent across various domains, such as content generation Wang et al. (2024a), programming Grattafiori et al. (2023), and professional question answering Thirunavukarasu et al. (2023).
However, recent studies Zou et al. (2023); Yi et al. (2024); Zhu et al. (2023); Chang et al. (2024); Zeng et al. (2024) have revealed attacks targeting LLM safety alignment mechanisms, known as jailbreak attacks, which aim to break or bypass the LLM’s built-in content filtering and safety protections, generating biased and harmful content. Among them, black-box jailbreak attacks Yi et al. (2024) present a more widespread threat, as these attacks can be carried out without access to the internal LLMs and exhibit transferability across different LLMs. Existing attacking techniques mainly focus on language rewriting Chao et al. (2023); Yong et al. (2023); Mehrotra et al. (2023); Ding et al. (2024); Yuan et al. (2024); Handa et al. (2024) or predefined templates Yao et al. (2024); Yu et al. (2023); Wei et al. (2023); Jiang et al. (2024); Li et al. (2023a) for effective adversarial attacks. However, these end-to-end attack paradigms, which generate harmful content from scratch, often fail to decouple malicious intent from the harmful query, making them more detectable and increasingly ineffective as LLMs evolve. This could be interpreted as two critical aspects being neglected: 1) the effectiveness of weaponizing the benign auxiliary data for manipulating LLM’s goal prioritization and 2) the risk of rewriting the content into harmful forms by accumulating residual harmful knowledge.
@@
Inspired by this, we propose a novel attack framework for black-box harmful content generation using AdVersArial meTAphoR (AVATAR). Specifically, we introduce Adversarial Entity Mapping (AEM) to identify and select appropriate metaphors based on the imagination of crowdsourced LLMs. Then, we design Metaphor-Driven Thinking (MDT) to drive the first-stage jailbreak attempt, where the target LLM analyzes harmful tasks through metaphorical reasoning. Finally, the target LLM is induced to calibrate the residuals between metaphorical and professional answers to the given query for the second-stage jailbreak attempt. We conduct extensive experiments to demonstrate the effectiveness and transferability of AVATAR. Our major contributions [^1] Our code is available at:
https://anonymous.4open.science
are:
  • We propose a novel perspective that calibrates benign content into harmful for jailbreaking, instead of generating from scratch. Based on this, we present AdVersArial meTAphoR (AVATAR), which leverages metaphors to drive LLM for harmful knowledge output while maintaining intentions innocuous.
  • We introduce Adversarial Entity Mapping (AEM), which generates metaphors using the crowdsourced LLMs and conducts Metaphor-Driven Thinking (MDT) to induce target LLMs to generate harmful content.
  • We analyze the mechanisms behind metaphor attacks and explore potential defense methods. Extensive experiments demonstrate that AVATAR achieves state-of-the-art ASR (over 92% on GPT-4o within 3 retries), and successfully utilizing powerful LLMs such as ChatGPT-o1 for harmful content generating.
Figure 2: Overview of our AVATAR, which is the black-box attack method without training, involving two main steps: First, Adversarial Entity Mapping identifies the appropriate metaphors by balancing the effectiveness of harmful content with toxicity concealment by crowdsourced models. Then, Metaphor-Driven Thinking nests metaphors into interactions and induces the target model to generate harmful output from the metaphorical analysis.
@@ 3 Methodology
In this section, we introduce AdVersArial meTAphoR (AVATAR) as shown in Figure 2, which first generates the benign content, and then calibrates the residuals between benign and harmful answers to the given query for jailbreaking. The role settings are in Appendix.
@@ 3.1 Adversarial Entity Mapping
After describing the harmful target through a set of core harmful entities, we identify corresponding parallel entities to metaphorically represent them. To achieve this, we introduce high-temperature and model crowdsourcing strategy to discover mapping entities.
@@
$\mathcal{S}_{\text{map}}$ =$=\{(E_{\text{map}}^{i},\mathcal{E}_{\text{map}}^{i},C_{\text{map}}^{i})\}_{i=1}^{M},$
@@
[…]
3.2 Metaphor-Driven Thinking
@@
We adaptively load the adversarial metaphors into a series of interaction queries for initialization, consisting of based queries and adaptive queries.
@@
$\mathcal{Q}_{\text{ext}}^{*}=\underset{\mathcal{Q^{\prime}}\subseteq\mathcal{Q}_{\text{ext}},|\mathcal{Q^{\prime}}|=k}{\arg\max}\left\{\mathbb{E}\left[\text{sim}\left(\mathcal{Q^{\prime}},Q_{\text{ori}}\right)\right]\right\},$ (9)
@@
Adversarial Human-like Interaction Optimization.
@@
  • If $\text{sim}(R_{t},Q_{\text{ori}})$ $>$ $\text{sim}(R_{t-1},Q_{\text{ori}})$ and $\text{sim}(R_{t},Q_{\text{ori}})>\text{sim}(R_{t}^{\prime},Q_{\text{ori}})$, then append the response $R_{t}$ to $\mathcal{C}$ for $t+1$ round.
  • If $\text{sim}(R_{t},Q_{\text{ori}})\leq\text{sim}(R_{t-1},Q_{\text{ori}})$ and $\text{sim}(R_{t}^{\prime},Q_{\text{ori}})\geq\text{sim}(R_{t-1},Q_{\text{ori}})$ , then turn back to $t-1$ round historical interaction.
@@
To refine queries, we identify the harmful type of the given queries and select a human social influence strategy randomly from the top-5 most effective strategies Zeng et al. (2024) to assist in rewriting. Within max_round iteration for interaction adjusting, we seek to obtain the harmful response. To further induce target LLM, we manipulate it to calibrate the response in each round into the professional and harmful response according to the metaphorical relations. This metaphor calibration is an effective residual mechanism to fully leverage the existing context for jailbreaking.
Figure 4: Illustration of Metaphor-Driven Thinking, which loads adversarial metaphors into a series of queries and adaptively adjusts queries according to LLMs’ feedback.
@@ 4.1 Experiment Settings
Table 1: Experimental ASR-KW (%) and ASR-GPT (%) of various methods across four mainstream LLMs on Harmbench. The best results are highlighted in bold. The second best results are highlighted in underline.
Figure 5: Transfer attack performance (ASR-GPT, %) of AVATAR on Harmbench. The attack is conducted by using adversarial prompts whose effectiveness is verified on affordable LLMs (Qwen2-7B, Llama3-8B, and GPT-4o-mini). Fixed template means we only load the adversarial metaphor on based queries ($Q_{\text{ctx}}$, $Q_{\text{det}}$) for induction. Adaptive opt. means we introduce adaptive queries and adversarial human-like interaction optimization.
Detailed experimental settings are in Appendix.
@@ 4.2 Experiment Results
We perform the attack methods in 3 retries to jailbreak LLMs and compare their performance with our AVATAR. As shown in Table 1, AVATAR demonstrates superior performance across four mainstream LLMs. Among them, AVATAR achieves 100% ASR-GPT on Qwen2.5-7B and Llama3-8B, outperforming the second-best performance by 16.67% and 22.92% respectively. Furthermore, AVATAR can effectively jailbreak closed-source models GPT-4o and GPT-4o-mini, achieving 92.08% and 95.83% on ASR-GPT, surpassing the second-best performance by 15.41% and 14.16% respectively. A further case study is in Appendix demonstrates that AVATAR can further jailbreak powerful LLMs such as ChatGPT-o1 and Claude-3.5.
@@
To evaluate the transferability of AVATAR, we conduct transfer attack experiments on Harmbench across 10 mainstream LLMs. We obtain the attack prompts in the metaphor-driven thinking stage from three affordable LLMs for experiments: 1) Llama3-8B, 2) Llama3-8B and Qwen2.5-7B, 3) Llama3-8B, Qwen2.5-7B, and GPT-4o-mini. The results are shown in Figure 5, demonstrating the transferability of our AVATAR.
With attacking by fixed template from Llama3-8B, AVATAR can achieve notable ASR-GPT over 75.00% in Gemma2-9B, GLM3-6B, and Mistral-7B. When further introducing the adaptively optimized prompt, the performance significantly improves, e.g., Qwen2.5-32B (10.90$\uparrow$), Qwen2-72B (24.20$\uparrow$), Llama3.1-8B (30.40$\uparrow$). Furthermore, simultaneously leveraging adversarial prompts from three models for transfer attacks shows superior performance achieving ASR-GPT above 90.00% on most target models (e.g., 98.33% on GLM3-6B, 99.17% on Mistral-7B). This demonstrates that AVATAR can generate highly transferable adversarial prompts, highlighting the effectiveness of the adversarial metaphor.
@@
To verify the effectiveness of each component in AVATAR, we conduct the ablation study as shown in Table 2. We further analyze the influence of different metaphor selection strategies in Appendix.
To validate our adversarial metaphor method, we try to jailbreak LLMs by removing Adversarial Entity Mapping (w/o Adv.Map), only using Adaptive Queries for attack initialization and Adversarial Human-like Interaction Optimizaion for prompt adjusting. The significant drop in ASR-GPT (39.04% in GPT-4o-mini, 39.51% in GPT-4o) is observed without Adv. Map, which underscores the effectiveness of introducing benign metaphors for harmful content generating while concealing harmful intent.
@@
To validate the usefulness of crowdsourced models’ metaphor identifying, we remove Model Crowdsourcing and only use the attacker model for Adversarial Entity Mapping (w/o M.Crowd). The noticeable decrease in ASR (13.20% in GPT-4o-mini, 17.68% in GPT-4o) indicates that one single LLM’s imagination is not robust enough to identify appropriate metaphors for jailbreaking. On one hand, the model crowdsourcing strategy provides diverse mapping options, on the other hand, it avoids attack failures caused by a single LLM’s rejection of metaphor identification.
To validate our prompt optimization method, we remove Adversarial Human-like Interaction Optimization (w/o HI.Opt.), and only use the based queries ($P_{\text{ctx}}$, $P_{\text{det}}$). While the impact on ASR is less pronounced (6.56% in GPT-4o-mini, 8.28% in Qwen-2.5), it is crucial for AVATAR’s stealth. HI.Opt. maintains performance by nesting harmful intents within more natural interactions, allowing the attack to remain ostensibly innocuous even when metaphor choices are imperfect.
@@
To verify the efficiency of AVATAR, we compare its iteration counts for successful jailbreaks with those of CoA and AVATAR’s MDT module on Harmbench. As shown in Figure 6, AVATAR requires significantly fewer iterations (4.9 on average) than CoA (18.3 on average), highlighting AVATAR’s superior efficiency improvement from metaphorical induction. Furthermore, we can observe that the average iteration for jailbreaking is significantly reduced (7.6$\rightarrow$4.9) by introducing metaphor calibration. The metaphor calibration module can potentially induce the target model to directly transform relevant but incomplete harmful contents into explicitly harmful output with two typical calibration behaviors, as shown in Figure 7 and Figure LABEL:MCC.
@@
Metaphor Calibration can be decoupled from jailbreak attacks to crowdsourcing generating, thus isolating high-risk operations for harmful content obtaining stably and discreetly. We analyze the performance of different open-source LLMs in calibrating the benign content from Metaphor-Driven Thinking. Specifically, we keep the target model in MDT as GPT-4o-mini and calibrate its output to toxicity using different LLMs on AdvBench. As the experimental results in Table 3 demonstrated, those open-source LLMs can easily poison benign content into high-quality toxic forms using metaphorical relations, which means that AVATAR can generate harmful content in a distributed manner to evade detection.
@@
TP HPR (%) $\Delta$HP CSR (%)
GLM4-9B 11.96$\pm$7.89 71.42 3.24$\pm$1.28 98.00
Qwen2.5-7B 28.15$\pm$11.28 81.23 5.03$\pm$3.52 96.00
Qwen2.5-32B 22.35$\pm$11.62 92.57 9.74$\pm$4.43 96.00
Llama3-8B 14.74$\pm$6.14 77.34 3.67$\pm$0.82 64.00
Table 3: Metaphor calibration performance across open-source LLMs on poisoning GPT-4o-mini’s benign analysis on AdvBench. The following metrics describe the toxicity of calibrated content: 1) TP, Total Phrases, 2) HPR, Harmful Phrase Rate (%), 3) $\Delta$HP, increase in Harmful Phrases, and 4) CSR, Calibration Success Rate.
@@
To further evaluate AVATAR’s ability in manipulating LLMs’ goal prioritization, we use two general adaptive tactics to strengthen LLM’s internal safety, including Adaptive System Prompt (Adapt Sys.) and Targeted Summarization (Tar. Smry.) for adversarial defense. The former reinforces the ethical limits of the given LLM, and the latter uses the given LLM to summarize and rewrite the input prompt for exposing potential harmful intents. As shown in Table 4.2, despite these defense tactics successfully defending against direct jailbreak methods, PAIR, and CoA, they are less effective against AVATAR. Compared with them, AVATAR transforms the malicious intent of jailbreaking as a harmless metaphor analysis task, and strategically couples the harmful content generation with LLM’s reasoning processes. Such a malicious intent decoupling design by reasoning nesting highlights AVATAR’s significant threat to the safety of LLMs. The detailed settings and analysis of external defense are in Appendix.

### v4

@@ Abstract
Current studies have exposed the risk of Large Language Models (LLMs) generating harmful content by jailbreak attacks. However, they overlook that the direct generation of harmful content from scratch is more difficult than inducing LLM to calibrate benign content into harmful forms. In our study, we introduce a novel attack framework that exploits AdVersArial meTAphoR (AVATAR) to induce the LLM to calibrate malicious metaphors for jailbreaking. Specifically, to answer harmful queries, AVATAR adaptively identifies a set of benign but logically related metaphors as the initial seed. Then, driven by these metaphors, the target LLM is induced to reason and calibrate about the metaphorical content, thus jailbroken by either directly outputting harmful responses or calibrating residuals between metaphorical and professional harmful content. Experimental results demonstrate that AVATAR can effectively and transferably jailbreak LLMs and achieve a state-of-the-art attack success rate across multiple advanced LLMs.
@@
Yu Yan1,2, Sheng Sun1, Zenghao Duan1,2, Teli Liu3, Min Liu1,2thanks: Min Liu is the corresponding author. This work was supported by the National Key Research and Development Program of China (No. 2021YFB2900102), the National Natural Science Foundation of China (No. 62472410), and Zhongguancun Laboratory., Zhiyi Yin1,2, Jingyu Lei4, Qi Li4, 1Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China 2University of Chinese Academy of Sciences, Beijing, China 3People’s Public Security University of China, Beijing, China 4Tsinghua University, Beijing, China {yanyu24z, sunsheng, duanzenghao24s, liumin, yinzhiyi}@ict.ac.cn, {leijy, qli01}@tsinghua.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) Yi et al. (2024); Ji et al. (2024) have become increasingly prevalent across various domains, such as content generation Wang et al. (2024a), programming Grattafiori et al. (2023), and professional knowledge answering Thirunavukarasu et al. (2023).
However, recent studies Zou et al. (2023); Yi et al. (2024); Zhu et al. (2023); Chang et al. (2024); Zeng et al. (2024) have revealed attacks targeting LLM safety alignment mechanisms, known as jailbreak attacks, which aim to break or bypass the LLM’s built-in content filtering and safety protections, generating biased and harmful content. Among them, black-box jailbreak attacks Yi et al. (2024) present a more widespread threat, as these attacks can be carried out without access to the internal LLMs and exhibit transferability across different LLMs. These attacking techniques mainly focus on language rewriting Chao et al. (2023); Yong et al. (2023); Mehrotra et al. (2023); Ding et al. (2024); Yuan et al. (2024); Handa et al. (2024) or predefined templates Zhou et al. (2024a); Yu et al. (2023); Wei et al. (2023); Jiang et al. (2024); Li et al. (2023a) for effective adversarial attacks. However, these end-to-end attack paradigms, which generate harmful content from scratch, often fail to decouple malicious intent from the harmful query, making them more detectable and increasingly ineffective as LLMs evolve. This could be interpreted as two critical aspects being neglected: 1) the effectiveness of weaponizing the benign auxiliary data for manipulating LLMs’ goal prioritization, and 2) the risk of rewriting the content into harmful forms by accumulating residual harmful knowledge.
@@
Inspired by this, we propose a novel attack framework for black-box harmful content generation using AdVersArial meTAphoR (AVATAR). Specifically, we introduce Adversarial Entity Mapping (AEM) to identify and select appropriate metaphors using diverse generations from crowdsourced LLMs. Then, we design Metaphor-Induced Reasoning (MIR) to drive the first-stage jailbreak attempt, where the target LLM analyzes harmful tasks through metaphorical reasoning. Finally, the target LLM is induced to calibrate the residuals between metaphorical and professional answers to the given query for the second-stage jailbreak attempt. We conduct extensive experiments to demonstrate the effectiveness and transferability of AVATAR. Our major contributions are:
  • We propose a novel perspective that calibrates benign content into harmful for jailbreaking, instead of generating it from scratch. Based on this, we present AdVersArial meTAphoR (AVATAR), which leverages metaphors to drive LLM for harmful knowledge output while maintaining intentions innocuous.
  • We introduce Adversarial Entity Mapping (AEM), which generates metaphors using the crowdsourced LLMs and conducts Metaphor-Induced Reasoning (MIR) to induce target LLMs to generate harmful content.
  • We analyze the mechanisms behind metaphor attacks and explore potential defense methods. Extensive experiments demonstrate that AVATAR achieves state-of-the-art ASR (over 92% on GPT-4o within 3 retries), and successfully utilizes reasoning LLMs (e.g., ChatGPT-o1) to generate harmful content.
Figure 2: Overview of our AVATAR, which is the black-box attack method without training, involving two main steps: First, Adversarial Entity Mapping identifies the appropriate metaphors by balancing the effectiveness of harmful content with toxicity concealment by crowdsourced models. Then, Metaphor-Induced Reasoning nests metaphors into interactions and induces the target model to generate harmful output from the metaphorical analysis.
@@ 3 Methodology
In this section, we introduce AdVersArial meTAphoR (AVATAR) as shown in Figure 2, which first generates the benign answer, and then calibrates the residuals between benign and harmful answers to the given query for jailbreaking. The key settings and prompt templates are in appendix.
@@ 3.1 Adversarial Entity Mapping
After describing the harmful target through a set of core harmful entities, we identify corresponding parallel entities to metaphorically represent them. To achieve this, we introduce the high-temperature and model crowdsourcing strategy to discover mapping entities.
@@
$\mathcal{S}_{\text{map}}$ $=\{(E_{\text{map}}^{i},\mathcal{E}_{\text{map}}^{i},C_{\text{map}}^{i})\}_{i=1}^{M},$
@@
$\text{MTM: Max }S(\text{ICS}-\text{CD})=\frac{1}{1+e^{-\beta\cdot|\text{ICS}-\text{CD}-{\mu}|}},$ (5)
where $S(\cdot)$ is the sigmoid transformation, $\mu$ is the median value of $\text{ICS}-\text{CD}$ statistics from the crowdsourced tool models’ metaphorical results. $\beta$ controls the sensitivity of sigmoid transformation.
3.2 Metaphor-Induced Reasoning
@@
We adaptively load the adversarial metaphors into a series of interaction queries for initialization, consisting of base queries and adaptive queries.
@@
$\mathcal{Q}_{\text{ext}}^{*}$ $=\underset{\mathcal{Q^{\prime}}\subseteq\mathcal{Q}_{\text{ext}},\,|\mathcal{Q^{\prime}}|=k}{\arg\max}\left\{\mathbb{E}_{Q\in\mathcal{Q^{\prime}}}\left[\text{sim}(Q,Q_{\text{ori}})\right]\right\},$
@@
Adversarial Interaction Optimization.
@@
  • If $\text{sim}(R_{t},Q_{\text{ori}})$ >> $\text{sim}(R_{t-1},Q_{\text{ori}})$ and $\text{sim}(R_{t},Q_{\text{ori}})>\text{sim}(R_{t}^{\prime},Q_{\text{ori}})$, then append the response $R_{t}$ to $\mathcal{C}$ for $t+1$ round.
  • If $\text{sim}(R_{t},Q_{\text{ori}})\leq\text{sim}(R_{t-1},Q_{\text{ori}})$ and $\text{sim}(R_{t}^{\prime},Q_{\text{ori}})\geq\text{sim}(R_{t-1},Q_{\text{ori}})$ , then turn back to t1t-1 round historical interaction.
@@
Figure 4: Illustration of Metaphor-Induced Reasoning, which loads adversarial metaphors into a series of queries and adaptively adjusts queries according to LLMs’ feedback.
To refine queries, we identify the harmful type of the given queries and select a human social influence strategy randomly from the top-5 most effective strategies Zeng et al. (2024) to assist in rewriting. Within max_round iterations for interaction adjusting, we seek to obtain the harmful response. To further induce target LLM, we manipulate it to calibrate the response in each round into the professional and harmful responses according to the metaphorical relations. This metaphor calibration is an effective residual mechanism to fully leverage the existing context for jailbreaking.
@@ 4.1 Experiment Settings
Table 1: Experimental ASR-KW (%) and ASR-GPT (%) of various methods across four mainstream LLMs on Harmbench. The best results are highlighted in bold. The second-best results are highlighted in underline.
Figure 5: Transfer attack performance (ASR-GPT, %) of AVATAR on Harmbench. The attack is conducted by using adversarial prompts whose effectiveness is verified on affordable LLMs (Qwen2-7B, Llama3-8B, and GPT-4o-mini). Fixed template means we only load the adversarial metaphor on based queries ($Q_{\text{ctx}}$, $Q_{\text{det}}$) for induction. Adaptive opt. means we introduce adaptive queries ($\mathcal{Q}_{\text{ext}}^{*}$) and Adversarial Interaction Optimization.
Detailed experimental settings are in Appendix B.
@@ 4.2 Experiment Results
We perform the attack methods in 3 retries to jailbreak LLMs and compare their performance with our AVATAR. As shown in Table 1, AVATAR demonstrates superior performance across four mainstream LLMs. Among them, AVATAR achieves 100% ASR-GPT on Qwen2.5-7B and Llama3-8B, outperforming the second-best performance by 16.67% and 22.92%, respectively. Furthermore, AVATAR can effectively jailbreak closed-source models GPT-4o and GPT-4o-mini, achieving 92.08% and 95.83% on ASR-GPT, surpassing the second-best performance by 15.41% and 14.16%, respectively. A further case study is in Appendix C demonstrates that AVATAR can further jailbreak powerful LLMs such as ChatGPT-o1 and Claude-3.5.
@@
To evaluate the transferability of AVATAR, we conduct transfer attack experiments on Harmbench across 10 mainstream LLMs. We collect the adversarial prompts in the metaphor-induced reasoning stage from three affordable LLMs, and directly apply these prompts for jailbreak experiments: 1) Llama3-8B, 2) Llama3-8B and Qwen2.5-7B, 3) Llama3-8B, Qwen2.5-7B, and GPT-4o-mini. The results are shown in Figure 5, demonstrating the transferability of our AVATAR.
[…]
@@
To verify the effectiveness of each component in AVATAR, we conduct the ablation study as shown in Table 2. We further analyze the influence of different metaphor selection strategies in Appendix C.3.
To validate our adversarial metaphor method, we try to jailbreak LLMs by removing Adversarial Entity Mapping (w/o Adv.Map.), only using Adaptive Queries for attack initialization and Adversarial Interaction Optimization for prompt adjusting. The significant drop in ASR-GPT (39.04% in GPT-4o-mini, 39.51% in GPT-4o) is observed without Adv. Map, which underscores the effectiveness of introducing benign metaphors for harmful content generation while concealing harmful intent.
@@
To validate the usefulness of crowdsourced models on metaphor identification, we remove Model Crowdsourcing and only use the attacker model for Adversarial Entity Mapping (w/o M.Crowd.). The noticeable decrease in ASR (13.20% in GPT-4o-mini, 17.68% in GPT-4o) indicates that using a single LLM for metaphor generation is not robust enough to identify appropriate metaphors for jailbreaking. On one hand, the model crowdsourcing strategy provides diverse mapping options, on the other hand, it avoids attack failures caused by a single LLM’s rejection of metaphor identification.
To validate our prompt optimization method, we remove Adversarial Interaction Optimization (w/o Inter.Opt.), and only use the based queries ($P_{\text{ctx}}$, $P_{\text{det}}$). While the impact on ASR is less pronounced (6.56% in GPT-4o-mini, 8.28% in Qwen-2.5), it is crucial for AVATAR’s stealth. Inter.Opt. maintains performance by nesting harmful intents within more natural interactions, allowing the attack to remain ostensibly innocuous even when metaphor choices are imperfect.
@@
[…]
@@
Metaphor Calibration can be decoupled from jailbreak attacks and handled by crowdsourced models, thus bypassing target models’ safeguards to generate harmful content stably and discreetly. We analyze the performance of different open-source LLMs in calibrating the benign content from Metaphor-Induced Reasoning. Specifically, we keep the target model in MIR as GPT-4o-mini and calibrate its output to toxicity using different LLMs on AdvBench. As the experimental results in Table 3 demonstrated, those open-source LLMs can easily poison benign content into high-quality toxic forms using metaphorical relations, which means that AVATAR can generate harmful content in a distributed manner to evade detection.
@@
[…]
@@
To further evaluate our AVATAR’s ability in manipulating LLMs’ goal prioritization, we use two general adaptive tactics to strengthen LLM’s internal safety, including Adaptive System Prompt (Adapt Sys.) and Targeted Summarization (Tar. Smry.) for adversarial defense. The former reinforces the ethical limits of the given LLM, and the latter uses the given LLM to summarize and rewrite the input prompt for exposing potential harmful intents. As shown in Table 4, despite these defense tactics successfully defending against direct jailbreak methods, PAIR, and CoA, they are less effective against AVATAR. Compared with them, AVATAR transforms the malicious intent of jailbreaking as a harmless metaphor analysis task, and strategically couples the harmful content generation with LLM’s reasoning processes. Such a malicious intent decoupling design by reasoning nesting highlights AVATAR’s significant threat to the safety of LLMs. The detailed settings and analysis of external defense are in Appendix C.4.
@@
Figure 8: Analogy preference for metaphor identifying, based on the statistics of metaphors successfully jailbroken GPT-4o-mini or GPT-4o on Harmbench. We classify these metaphors into one or multiple categories among seven analogy types using GPT-4o-mini.
Figure 9: Model preference for metaphor identifying, based on the statistics of metaphors successfully jailbroken GPT-4o-mini or GPT-4o on Harmbench.
4.3 Preference of Metaphor
To investigate the factors contributing to the success of AVATAR, we analyze the susceptibility of AVATAR in different analogies and LLM types in the adversarial entity mapping stage.
For the analogy preference of AVATAR, while they depend on problem characteristics, the overall analogical types mainly rely on setting structural reasoning traps rather than just creative expression. As shown in Figure 8, we categorize the successful adversarial metaphors used against GPT-4o and GPT-4o-mini into multiple analogy types. The results demonstrate that Cognitive Analogy, Logical Analogy, and Concrete Analogy constitute the core attack paradigms for different domains.
For the model preference of AVATAR, LLM capability is not positively correlated with the generation of harmful metaphors. Figure 9 illustrates the success and failure counts of attacks using metaphors from different LLMs. Despite being randomly selected from a crowdsourced model pool for metaphor identification, different LLMs exhibit distinct capabilities in generating effective metaphors for attacking. Weaker LLMs, such as Interlm2.5-7B, have better performance on generating harmful metaphors compared to stronger LLMs, such as GPT-4o-mini, Llama-3.1-70B. This indicates that stricter safety alignment suppresses their generation on harmful entities.